OpenStudy (anonymous):

Hi all! I have a question :) "An open rectangular box with volume 2m^3 has a square base. Express the surface area of the box as a function of the length of a side of the base"

OpenStudy (anonymous):

:/

OpenStudy (lena772):

Area of base is x^2 Volume = base * height 2 = x^2 * h h = 2/ x^2 4(2 / x^2)(x) + x^2 = SA SA = 8/x + x^2

OpenStudy (anonymous):

how do you figure out the height? That was my problem